OpenStudy (swaqqout_kid1):

Which value for x makes the sentence true? 2x - 1 = 9 x = 10 x = 2 x = 12 x = 5 @EmmaMink

OpenStudy (emmamink):

Let's see how to solve this.....substitute in each value as x, do the operations and find out which one is true! I'll help you out!

OpenStudy (kamibug):

Add 1 to both sides then divide both sides by 2.

OpenStudy (igreen):

2x - 1 = 9 Add 1 to both sides. 2x = 9 + 1 2x = 10 Now, can you divide 10 / 2?

OpenStudy (anonymous):

5

OpenStudy (emmamink):

Let's start with adding one and then dividing each side by two. This is called the addition property of equality and the multiplication property of equality. We now have x = 5!

OpenStudy (igreen):

Yes, that is your answer.

OpenStudy (anonymous):

Emmamink, you are wrong

OpenStudy (emmamink):

Please, if appreciated medal.

OpenStudy (emmamink):

How? The answer is 5!

OpenStudy (emmamink):

2 times 5 minus 1 is 9...is it not?

Nnesha (nnesha):

if you don't how to solve for x so simply plug x values into the equation  both side not equal so cancel out A